# {strip}[Uncaptioned image]

### v1

@@ 4.2. Ablation study
The ablation studies are conducted in two parts. First, we investigate the effects of incorporating text and image embeddings at different positions within the cross-attention. Second, we evaluate the impact of fine-tuning different modules during the third stage of training.
@@ 4.2.1. Inject position of image and text embedding
In Sec. 3.4, we introduce the Image Representation and Textual Semantics Decoupling framework, where image and text embeddings are injected as cross-attention at different position of the denoising network. We presents an ablation study on the specific injection positions of both embeddings in Sec. 4.2.1.
The locations of image and text embedding injection within LoopAnimation are detailed in Fig. 2, and determined through a combination of subjective and objective evaluations on a test set of 99 images. Subjective evaluations involved seven participants who scored the generated animations from low to high (1 being the lowest and 3 being the highest) based on two aspects: subject consistency (${Q}_{S}$) and motion quality (${Q}_{M}$). We also calculated the Loop-C objective metric to assess loopability. The results are presented in Table 4. The experiment group achieving the highest ${Q}_{S}$ score utilized only image embeddings. However, observations revealed minimal motion dynamics, leading to a low ${Q}_{M}$ score and a high Loop-C metric due to the restricted motion range.
The subjective visual results of the experiments are presented in Fig. 4. Fig. 4 showcases five sets of experimental outcomes, each corresponding to the configurations detailed in Table 4. The structure employed within LoopAnimate is highlighted in bold within Fig. 4. In the second row of experiments, the absence of text embedding injection into the middle block layer results in minimal motion dynamics. Additionally, the generated videos lack specific movements guided by the text prompts, which is a prevalent observation within this experimental group. This aligns with the findings of  (basu2023localizing, ) regarding the critical role of cross-attention within the middle block for motion generation in text-to-image tasks. Therefore, we conclude that incorporating text embedding within the middle block is equally crucial for video generation tasks. The third row presents experiments where only text embedding information is injected into all layers. While motion dynamics are present, the lack of input image feature-level information leads to object disintegration and low-quality video generation. Conversely, the fourth row showcases experiments with exclusive image information injection into all layers, resulting in a near-complete loss of motion dynamics. This phenomenon is consistently observed in other unreported results from this group, leading to the conclusion that excessive image embedding injection can negatively impact the motion dynamics of the generated output.
@@ 5. Discussion and conclusion
This paper proposes LoopAnimate, a novel video generation method capable of producing videos with seamless loops between the first and last frames. Notably, LoopAnimate extends the one-shot video generation length of UNet architectures to 35 frames, a significant advancement in the field. We conducted extensive experiments on a custom-built dataset of subject images, comparing LoopAnimate with existing image-to-video methods. Both objective and subjective evaluation results demonstrate that LoopAnimate outperforms current approaches in terms of subject preservation, inter-frame continuity, and motion quality, validating its effectiveness.

### v2

@@ 4.2. Ablation study
The ablation studies are conducted in two parts. First, we investigate the effects of incorporating image and text embeddings at different positions within the cross-attention. Second, we evaluate the impact of fine-tuning different modules during the third stage of training.
@@ 4.2.1. Inject position of image and text embedding
In Sec. 3.4, we introduce the Multi-level Image representation and Textual semantics Decoupling Framework, where image and text embeddings are injected as cross-attention at different position of the denoising network. We presents an ablation study on the specific injection positions of both embeddings in Sec. 4.2.1.
The locations of image and text embedding injection within LoopAnimation are detailed in Figure. 2, and determined through a combination of subjective and objective evaluations on a test set of 99 images. Subjective evaluations involved seven participants who scored the generated animations from low to high (1 being the lowest and 3 being the highest) based on two aspects: subject consistency (${Q}_{S}$) and motion quality (${Q}_{M}$). We also calculated the Loop-C objective metric to assess loopability. The results are presented in Table 4. The experiment group achieving the highest ${Q}_{S}$ score utilized only image embeddings. However, observations revealed minimal motion dynamics, leading to a low ${Q}_{M}$ score and a high Loop-C metric due to the restricted motion range.
The subjective visual results of the experiments are presented in Figure. 4. Figure. 4 showcases five sets of experimental outcomes, each corresponding to the configurations detailed in Table 4. The structure employed within LoopAnimate is highlighted in bold within Figure. 4. In the second row of experiments, the absence of text embedding injection into the middle block layer results in minimal motion dynamics. Additionally, the generated videos lack specific movements guided by the text prompts, which is a prevalent observation within this experimental group. This aligns with the findings of  (basu2023localizing, ) regarding the critical role of cross-attention within the middle block for motion generation in text-to-image tasks. Therefore, we conclude that incorporating text embedding within the middle block is equally crucial for video generation tasks. The third row presents experiments where only text embedding information is injected into all layers. While motion dynamics are present, the lack of input image feature-level information leads to object disintegration and low-quality video generation. Conversely, the fourth row showcases experiments with exclusive image information injection into all layers, resulting in a near-complete loss of motion dynamics. This phenomenon is consistently observed in other unreported results from this group, leading to the conclusion that excessive image embedding injection can negatively impact the motion dynamics of the generated output.
@@ 5. Discussion and conclusion
This paper proposes LoopAnimate, a novel video generation method capable of producing videos with seamless loops between the first and last frames. Notably, LoopAnimate extends the one-shot video generation length to 35 frames, which is a significant advancement in the field. We conducted extensive experiments on a in-house dataset of subject images, comparing LoopAnimate with existing image-to-video methods. Both objective and subjective evaluation results demonstrate that LoopAnimate outperforms current approaches in terms of subject preservation, inter-frame continuity, and motion quality, validating its effectiveness.